# Practical hybrid decoding scheme for parity-encoded spin systems

### v2

@@ IV Experimental demonstration
The performance of the BF algorithm was investigated under the assumption of i.i.d. noise, the simplest model of noisy readout in QA. Pastawski and Preskill previously studied this model using the BP algorithm Pastawski and Preskill (2016). Note that the BF decoding in Eq.(36) depends only on the weight-3 syndromes, which treat all variables (VN and CN) symmetrically Pastawski and Preskill (2016). Furthermore, if we note that the LDPC codes are linear codes and assume the AWGN channel is symmetric, i.e. $P(y_{i}|z_{i})=P(-y_{i}|-z_{i})$ holds, the probability of successful decoding is independent of the input code-state $\boldsymbol{z}$ Richardson and Urbanke (2001). In this case, without loss of generality, the performance of decoding can be analyzed using the assumption that all-one code-state $\boldsymbol{z}=\left(+1,\ldots,+1\right)$ has been transmitted Richardson and Urbanke (2001); Vuffray (2014). In physics, this assumption corresponds to choosing the ferromagnetic gauge and identifying the current decision $\boldsymbol{r}$ with an error pattern $\boldsymbol{e}$. Thus, we assume the input is all-one code-state, that is, $\hat{\boldsymbol{z}}=\hat{\boldsymbol{1}}_{K\times K}$. All the demonstrations were performed using the Mathematica® Ver.14 platform on the Windows 11 operating system. We generated 5000 symmetric matrices $\hat{\boldsymbol{r}}=\hat{\boldsymbol{e}}$ with unit diagonal elements and other elements randomly assigned 1-1 with probability $\varepsilon<\tfrac{1}{2}$ and +1+1 otherwise. After $n=5$ iterations, it was checked whether the errors were eliminated. Figure 6(a) shows the performance of the BF algorithm, plotting the probability of decoding failure as a function of the number $K$ of logical spins ranging from $2$ to $40$ (associated SLHZ model consists of $\tbinom{K}{2}$ physical spins) for seven values of common bit error rate $\varepsilon$ $\left(=0.05,0.07,0.1,0.15,0.2,0.3,0.4\right)$. The MATLAB code implementing the described BF decoder and producing the figure is included as Supplemental Material *[SeeSupplementalMaterialat][forMATLABcodeimplementationofdescribedBFdecoding.]matlabcode. The failure probability falls steeply as $N$ increases if $\varepsilon$ is not too close to the threshold value $1/2$. A similar performance calculation was reproduced for the decoding using the BP algorithm after being iterated five times, as shown in Fig.6(b) Pastawski and Preskill (2016). Comparing these figures shows that the performance of the BF algorithm is comparable to that of the BP algorithm. Note that the BP algorithm updates all marginal probabilities $P\left(x_{i}|\boldsymbol{y}\right)$ associated with the $\tbinom{K}{2}$ spins sequentially by passing a real-valued message between the associated VNs and CNs per a single iteration. In contrast, in the BF algorithm, all the $\tbinom{K}{2}$ variables $r_{ij}$ in the matrix $\hat{\boldsymbol{r}}$ are updated in parallel per a single iteration. Thus, each spin variable was updated five times in these algorithms.
@@
where $J_{ij}\in\mathbb{R}$ is a coupling constant which is identified with the channel observation $y_{ij}\in\mathbb{R}$ in the AWGN channel model. Parameters $\left\{\beta,\gamma\right\}$ are independent annealing parameters to be adjusted. If the weight-4 syndrome is defined by
@@
Here, we must draw attention to coupling constants $\{J_{ij}\}$. Recall that the primary purpose of this study is to propose and demonstrate a practical method for readout decoding of the SLHZ-based QA devices. In this context, $\{J_{ij}\}$ is given by the problem we want to tackle. For example, they are given by an externally defined COP. In other cases, we can consider minimization of the Hamiltonian (50) as a decoding of the observation $\boldsymbol{J}=(J_{12},\ldots,J_{K-1\,K})\in\mathbb{R}^{\tbinom{K}{2}}$ in the AWGN channel model. To avoid duplicative consideration in analyzing the potential of the BF algorithm for readout decoding of the SLHZ model, we disregarded the reliability information for each spin readout of the SLHZ model itself, even if it were available in the actual QA device. Thus, the following results can be considered valid if readout errors are solely due to leakage errors arising from dynamical and thermal excitations during the sampling process, while neglecting other errors, such as measurement errors.
As a starting point, we provide typical examples demonstrating that both our BF and BP algorithms successfully eliminate errors in the readout of the SLHZ model generated by a classical MCMC sampler. In this example, we simulated readouts $\hat{\boldsymbol{r}}=\hat{\boldsymbol{e}}\circ\hat{\boldsymbol{z}}\in\{\pm 1\}^{K\times K}$ of the SLHZ system associated with a spin glass problem for $K=14$ ($K_{14}$) as a toy model. We generated a set of $12$ logical random instances on complete graphs $K_{14}$ with couplings $J_{ij}\in[-\tfrac{1}{4},\tfrac{1}{4}]$ chosen uniformly at random and all logical local fields set to zero, where the code-state $\hat{\boldsymbol{z}}$ for a given $\boldsymbol{J}$ was precomputed by brute force. The leftmost matrix plot in Fig.8 visualizes a typical example of the error matrix $\hat{\boldsymbol{e}}$ associated with a sampled readout $\hat{\boldsymbol{r}}$ of the SLHZ system. In this example, the error distribution was quite different from the expected for an i.i.d. noise model in Fig.7. Fig.8 shows the results where $\hat{\boldsymbol{e}}$ was decoded by the BF algorithm (upper) and the BP algorithm (lower) when the number of iterations was limited to 5 times. These results suggest that our BF algorithm, as well as the BP algorithm, can eliminate errors in the stochastically sampled readouts of the SLHZ system, which does not follow the i.i.d. noise model. Here, although we can not know the error rate for each spin because the correct ground state is generally unknown a priori, an initial error rate of 0.25 was assumed for every spin when running the BP algorithm.
@@
The performance of the two methods differs significantly. The columns (a) and (b) in Fig.9 indicate landscapes of the success probability for (a) the MCMC decoding and (b) the MCMC-BF hybrid decoding, respectively. In these figures, the top two figures are the success probability for finding the error-free code-state $\hat{\boldsymbol{z}}$, and the bottom two figures are the probability for finding any code-state $\hat{\boldsymbol{x}}=\boldsymbol{X}^{T}\boldsymbol{X}$, where $\boldsymbol{X}\in\{\pm 1\}^{K}$ is any logical state. It is very important to note that the sizes of the sample $\{\hat{\boldsymbol{x}}\}$ in the MCMC decoding and the decoded sample $\{\hat{\boldsymbol{r}}\}$ in the MCMC-BF hybrid decoding differ significantly; they were $1200\tbinom{K}{2}$ in the MCMC decoding and $4\tbinom{K}{2}$ in the MCMC-BF hybrid decoding. This indicates that the number of iterations required for MCMC sampling was 300 times smaller for the MCMC-BF hybrid decoding than for the MCMC decoding. Therefore, the MCMC-BF hybrid decoding provides a better trade-off between error performance and decoding complexity if the decoding complexity of the BF decoding is negligible. We will revisit the validity of this assumption later.
@@
Let us note two arrows $A$ and $B$ shown in Fig.9. They correspond to the annealing parameter sets $A=\{\beta_{A},\gamma_{A}\}$ and $B=\{\beta_{B},\gamma_{B}\}$ for which the success probability for decoding is maximized in the MCMC decoding (A) and the MCMC-BF hybrid decoding (B), respectively. The figure clearly indicates that the parameter sets suitable for use differ between the respective decoding methods. The left and right in Fig. 10 are the matrix plots showing the average error matrix $\langle\hat{\boldsymbol{e}}\rangle=\langle\hat{\boldsymbol{x}}\rangle\circ\hat{\boldsymbol{z}}$ over the set of samples $\{\hat{\boldsymbol{x}}\}$ obtained from the MCMC sampler under the parameter sets $A$ and $B$, respectively. The marginal probability $P_{c}$ that $\hat{x}_{ij}$ gives the correct value, i.e., $\hat{x}_{ij}=\hat{z}_{ij}$, is given by $P_{c}=\tfrac{1+\langle{\hat{e}_{ij}}\rangle}{2}$ . So, a negative entry in $\langle\hat{\boldsymbol{e}}\rangle$ indicates it is likely that there is an error in the corresponding entry in $\hat{\boldsymbol{x}}$. We can see that $\langle\hat{\boldsymbol{e}}\rangle$ depends on the chosen parameter sets. The average error matrix $\langle\hat{\boldsymbol{e}}\rangle$ in Fig.10(b) is consistent with the error matrix $\hat{\boldsymbol{e}}$ in Fig.8, because they were sampled with the same parameter set $B$. Note that it is highly likely to fail sampling a code-state by MCMC sampler for the set $B$ (see the lower left plot in Fig.9), while it is highly likely to succeed sampling a code-state by the MCMC sampler for the set $A$. This implies that the state $\hat{\boldsymbol{x}}$ sampled in the first stage of the MCMC-BF decoding is not a code-state, although the state $\hat{\boldsymbol{x}}$ sampled in the MCMC decoding is a code-state. This is quite reasonable because the subsequent BF decoding can correct errors only if $\hat{\boldsymbol{x}}$ sampled at the first stage is a leakage state in the MCMC-BF hybrid decoding. In contrast, the sampled state must be a code-state if it is the correct state in the MCMC decoding. This is why optimal parameter sets $\{\beta_{opt},\gamma_{opt}\}$ differ between the MCMC decoding and MCMC-BF hybrid decoding. It is important to note that this is a consequence of our BF decoding algorithm and is independent of the sampling algorithm in the first stage. Our results suggest that optimal annealing parameters should be carefully chosen when applying our BF decoding for readouts of QA as they may differ from the optimal parameters for QA when used alone.
Figure 9: Landscapes of the probability distribution for successful decoding. They are plotted as functions of the annealing parameters $\{\beta,\gamma\}$ in Eq.(50). The left and right columns are the results of (a) MCMC decoding and (b) MCMC-BF hybrid decoding, respectively. The target states in the upper and lower rows of these figures are different: they are the correct code-state $\hat{\boldsymbol{z}}$ for the upper row and any code-states for the lower row.
@@ V.1 Relationship between MCMC-BF hybrid algorithm and other known BF algorithms
Both the MCMC and BF decoding algorithms are hard-decision algorithms, in which only spin variables are of concern. The MCMC-BF decoding algorithm is considered to incorporate MCMC decoding as a preprocessing step for the BF decoding algorithm. We discuss the limitations of BF decoding and how they were resolved in the existing extended BF algorithms.
@@
Many researchers have developed sophisticated versions of the weighted BF (WBF) to incorporate reliability information of the hard decisions into the BF algorithm. See details in Refs.Khoa Le Trung (2017); Kennedy Masunda (2017) , and references therein. Alternatively, there is another approach to incorporate soft information. For example, Wadayama et al. proposed the Gradient Descent Bit Flipping (GDBF) algorithm Wadayama et al. (2010). They are based on modifications to the inversion function that may improve decoding performance, though this comes at the cost of increased decoding complexity. For example, the inversion functions employed in the BF, WBF, and GDBF algorithms are formally written as
@@
respectively. Here, $\boldsymbol{x}=(x_{1},\ldots,x_{N_{v}})\in\{\pm 1\}^{N_{v}}$ is the current decision for the spin variables (VNs), $\beta$ is a positive real parameter to be adjusted, and
@@
is an $i$th syndrome (CN) for the decision $\boldsymbol{x}$ in the spin representation, and $J_{k}$ is identified with channel observation $y_{k}$ in the AWGN channel model. Here, the set $N(i)=\left\{j:H_{ij}=1\right\}$ is the VNs adjacent to an $i$th CN $\left(1\leq i\leq N_{c}\right)$ and the set $M(j)=\left\{i:H_{ij}=1\right\}$ is the CNs adjacent to a $j$th VN $\left(1\leq j\leq N_{v}\right)$. Please refer to Fig.2 for the definition of these sets. Note that as long as $s_{i}\left(\boldsymbol{x}\right)$ is a weight-3 syndrome, $\varDelta_{k}^{\mathrm{(BF)}}(\boldsymbol{x})$ is symmetric for the permutation of the elements of $\boldsymbol{x}$. In other words, $\varDelta_{k}^{\mathrm{(WBF)}}(\boldsymbol{x})$ is invariant under exchanging $i\longleftrightarrow j$ for any pair of indices of VN other than $k$, justifying the all-one code-state assumption. In contrast, weight-4 syndrome is not symmetric. Similarly, because $\varDelta_{k}^{\mathrm{(WBF)}}(\boldsymbol{x})$ involves weight $w_{k}$ in the second term and $\varDelta_{k}^{\mathrm{(GDBF)}}(\boldsymbol{x})$ involves $J_{k}$ in the first term, both depend on $\boldsymbol{J}$. Thus, $\varDelta_{k}^{\mathrm{(WBF)}}(\boldsymbol{x})$ and $\varDelta_{k}^{\mathrm{(GDBF)}}(\boldsymbol{x})$ are not symmetric even if $s_{i}(\boldsymbol{x})$ is a weight-3 syndrome. It is interesting to note that the inversion functions in Eq.(52)-(54) can be formally derived from the following Hamiltonians:
@@
where $X=\mathrm{BF}$, $\mathrm{WBF}$, or $\mathrm{GDBF}$. It follows that if $\varDelta_{k}^{(X)}(\boldsymbol{x})<0$, flipping the $k$th spin reduces the total energy of the spins. Therefore, the BF, WBF, and GDBF decoding algorithms can be considered deterministic algorithms that determine the most suitable spins to be flipped to reduce the energy $H_{k}^{(X)}(\boldsymbol{x})$ based on the inverse function $\varDelta_{k}^{(X)}(\boldsymbol{x})$. In contrast to the BF algorithm, the BP algorithm essentially considers soft information, as it calculates a consistent marginal probability $P\left(x_{i}|\boldsymbol{y}\right)$ by exchanging real-valued messages between VNs and CNs. On the other hand, the inversion function and the associated Hamiltonian for the MCMC decoding are formally written by
@@
In addition, MCMC decoding in the first stage also solves another limitation of the BF algorithm in the second stage. The MCMC sampling is stochastic in contrast to our deterministic BF algorithm and gradient descent algorithm used in the GDBF algorithm. This stochasticity introduces randomness into the spin-flip selection. It allows spins to flip even when $\varDelta_{k}^{\mathrm{(MCMC)}}(\boldsymbol{x})>0$, which provides an escape from spurious local minima and makes it more likely to arrive at the neighborhood of the global minimum of $H^{\mathrm{(MCMC)}}(\boldsymbol{x})$. In this way, the stochastic spin-flip selection helps find the global minimum of $H^{\mathrm{(MCMC)}}(\boldsymbol{x})$. A similar stochasticity can be incorporated directly into the GDBF algorithm, either by adding a noise term to the inversion function $\varDelta_{k}^{\mathrm{(GDBF)}}(\boldsymbol{x})$ (Noisy GDBF Sundararajan et al. (2014)) or by taking stochastic spin-flip selection into account in the algorithm (Probabilistic GDBF Rasheed et al. (2014)). Briefly speaking, our MCMC-BF hybrid decoding algorithm can be considered an alternative to the BP algorithm and other existing BF decoding algorithms. Our simulation suggests that it is important to properly control the fluctuations of $\boldsymbol{x}$ during MCMC sampling by adjusting the two annealing parameters $\left\{\beta,\gamma\right\}$ to optimize MCMC-BF decoding. It also suggests that the contribution of the penalty term to the Hamiltonian must be small enough not to force $\boldsymbol{x}$ into a valid code-state. Controlling annealing parameters is important for switching between the two decoding modes. Fig.11 shows schematic diagrams illustrating two decoding modes: (a) MCMC decoding and (b) MCMC-BF hybrid decoding, respectively. Decoding is formulated as a constrained COP. The large and small ellipses indicate the search space and the code-state space that minimizes the penalty term of the Hamiltonian. In MCMC decoding, the MCMC samples both the code-states (feasible solution) and leakage states (infeasible solution) (Fig.11(a)). In some occasions, the correct code-state (solution) may be sampled. In MCMC-BF hybrid decoding, on the other hand, the first-stage MCMC sampler samples only the leakage state, and the second-stage BF decoding maps the leakage state to the correct code-state occasionally (Fig.11(b)). Our simulations showed that the MCMC-BF hybrid decoding (Fig.11(b)) was over two orders of magnitude more efficient than the MCMC decoding (Fig.11(a)) if we ignore the decoding cost of the post-readout BF decoding and compare performance simply by the number of required MCMC samples needed to obtain at least one correct code-state.

### v3

@@ IV Experimental demonstration
The performance of the BF algorithm was investigated under the assumption of i.i.d. noise, the simplest model of noisy readout in QA. Pastawski and Preskill previously studied this model using the BP algorithm [20]. Note that the BF decoding in Eq.(36) depends only on the weight-3 syndromes, which treat all variables (VN and CN) symmetrically [20]. Furthermore, if we note that the LDPC codes are linear codes and assume the AWGN channel is symmetric, i.e. $P(y_{i}|z_{i})=P(-y_{i}|-z_{i})$ holds, the probability of successful decoding is independent of the input code-state $\boldsymbol{z}$ [31]. In this case, without loss of generality, the performance of decoding can be analyzed using the assumption that all-one code-state $\boldsymbol{z}=\left(+1,\ldots,+1\right)$ has been transmitted [31, 32]. In physics, this assumption corresponds to choosing the ferromagnetic gauge and identifying the current decision $\boldsymbol{r}$ with an error pattern $\boldsymbol{e}$. Thus, we assume the input is all-one code-state, that is, $\hat{\boldsymbol{z}}=\hat{\boldsymbol{1}}_{K\times K}$. All the demonstrations were performed using the Mathematica® Ver.14 platform on the Windows 11 operating system. We generated 5000 symmetric matrices $\hat{\boldsymbol{r}}=\hat{\boldsymbol{e}}$ with unit diagonal elements and other elements randomly assigned 1-1 with probability $\varepsilon<\tfrac{1}{2}$ and +1+1 otherwise. After $n=5$ iterations, it was checked whether the errors were eliminated. Figure 6(a) shows the performance of the BF algorithm, plotting the probability of decoding failure as a function of the number $K$ of logical spins ranging from $2$ to $40$ (associated SLHZ model consists of $\tbinom{K}{2}$ physical spins) for seven values of common bit error rate $\varepsilon$ $\left(=0.05,0.07,0.1,0.15,0.2,0.3,0.4\right)$. The MATLAB code implementing the described BF decoder and producing the figure is included as Supplemental Material [*[SeeSupplementalMaterialat][forMATLABcodeimplementationofdescribedBFdecoding.]matlabcode]. The failure probability falls steeply as $N$ increases if $\varepsilon$ is not too close to the threshold value $1/2$. A similar performance calculation was reproduced for the decoding using the BP algorithm after being iterated five times, as shown in Fig.6(b) [20]. Comparing these figures shows that the performance of the BF algorithm is comparable to that of the BP algorithm. Note that the BP algorithm updates all marginal probabilities $P\left(x_{i}|\boldsymbol{y}\right)$ associated with the $\tbinom{K}{2}$ spins sequentially by passing a real-valued message between the associated VNs and CNs per a single iteration. In contrast, in the BF algorithm, all the $\tbinom{K}{2}$ variables $r_{ij}$ in the matrix $\hat{\boldsymbol{r}}$ are updated in parallel per a single iteration. Thus, each spin variable was updated five times in these algorithms.
@@
where $J_{ij}\in\mathbb{R}$ is a coupling constant which is identified with the channel observation $y_{ij}\in\mathbb{R}$ in the AWGN channel model. The first term in Eq. (50) corresponds to the distance metric between a codeword $\boldsymbol{x}$ and the received information $\boldsymbol{J}$, while the second term corresponds to the constraint satisfaction metric [35]. Parameters $\left\{\beta,\gamma\right\}$ are the Lagrange weights for these two terms.
If the weight-4 syndrome is defined by
@@
Here, we must draw attention to coupling constants $\{J_{ij}\}$. Recall that the primary purpose of this study is to propose and demonstrate a practical method for readout decoding of the SLHZ-based QA devices. In this context, $\{J_{ij}\}$ is given by the problem we want to tackle. For example, they are given by an externally defined COP. In other cases, we can consider minimization of the Hamiltonian (50) as a decoding of the observation $\boldsymbol{J}=(J_{12},\ldots,J_{K-1\,K})\in\mathbb{R}^{\tbinom{K}{2}}$ in the AWGN channel model. To avoid duplicative consideration in analyzing the potential of the BF algorithm for readout decoding of the SLHZ model, we disregarded the reliability information for each spin readout of the SLHZ model itself, even if it were available in the actual QA device. Thus, the following results can be considered valid if readout errors are solely due to leakage errors arising from dynamical and thermal excitations during the sampling process, while neglecting other errors, such as measurement errors. In the Appendix B, we present a supplementary result when i.i.d. noise is introduced during the readout stage. This demonstrates that the above simplified approach also exhibits reasonable robustness against i.i.d. readout errors.
@@
As a starting point, we provide typical examples demonstrating that both our BF and BP algorithms successfully eliminate errors in the readout of the SLHZ model generated by a classical MCMC sampler. In this example, we simulated readouts $\hat{\boldsymbol{r}}=\hat{\boldsymbol{e}}\circ\hat{\boldsymbol{z}}\in\{\pm 1\}^{K\times K}$ of the SLHZ system associated with a spin glass problem for $K=14$ ($K_{14}$) as a toy model. We generated a set of $12$ logical random instances on complete graphs $K_{14}$ with couplings $J_{ij}\in[-\tfrac{1}{4},\tfrac{1}{4}]$ chosen uniformly at random and all logical local fields set to zero, where the code-state $\hat{\boldsymbol{z}}$ for a given $\boldsymbol{J}$ was precomputed by brute force. The leftmost matrix plot in Fig.8 visualizes a typical example of the error matrix $\hat{\boldsymbol{e}}$ associated with a sampled readout $\hat{\boldsymbol{r}}$ of the SLHZ system. In this example, the error distribution was quite different from the expected for an i.i.d. noise model in Fig.7. Fig.8 shows the results where $\hat{\boldsymbol{e}}$ was decoded by the BF algorithm (upper) and the BP algorithm (lower) when the number of iterations was limited to 5 rounds. These results suggest that our BF algorithm, as well as the BP algorithm, can eliminate errors in the stochastically sampled readouts of the SLHZ system, which does not follow the i.i.d. noise model. Here, although we can not know the error rate for each spin because the correct ground state is generally unknown a priori, an initial error rate of 0.25 was assumed for every spin when running the BP algorithm.
Figure 9: Landscapes of the probability distribution for successful decoding. They are plotted as functions of the Lagrange weights $\{\beta,\gamma\}$ in Eq.(50). The left and right columns are the results of (a) MCMC decoding and (b) MCMC-BF hybrid decoding, respectively. The target states in the upper and lower rows of these figures are different: they are the correct code-state $\hat{\boldsymbol{z}}$ for the upper row and any code-states for the lower row.
@@
The performance of the two methods differs significantly. The columns (a) and (b) in Fig.9 indicate landscapes of the success probability for (a) the MCMC decoding and (b) the MCMC-BF hybrid decoding, respectively. In these figures, the top two figures are the success probability for finding the error-free code-state $\hat{\boldsymbol{z}}$, and the bottom two figures are the probability for finding any code-state $\hat{\boldsymbol{x}}=\boldsymbol{X}^{T}\boldsymbol{X}$, where $\boldsymbol{X}\in\{\pm 1\}^{K}$ is any logical state. It is very important to note that the sizes of the sample $\{\hat{\boldsymbol{x}}\}$ for the MCMC decoding and the decoded sample $\{\hat{\boldsymbol{r}}\}$ for the MCMC-BF hybrid decoding differ significantly; they were $1200\tbinom{K}{2}$ in the MCMC decoding and $4\tbinom{K}{2}$ in the MCMC-BF hybrid decoding. This indicates that the number of iterations required for MCMC sampling was 300 times smaller for the MCMC-BF hybrid decoding than for the MCMC decoding. Therefore, the MCMC-BF hybrid decoding provides a better trade-off between error performance and decoding complexity if the decoding complexity of the BF decoding is negligible. We will revisit the validity of this assumption later.
@@
Let us note two arrows $A$ and $B$ shown in Fig.9. They correspond to the sets of Lagrange weights $A=\{\beta_{A},\gamma_{A}\}$ and $B=\{\beta_{B},\gamma_{B}\}$ for which the success probability for decoding is maximized in the MCMC decoding (A) and the MCMC-BF hybrid decoding (B), respectively. The figure clearly indicates that the parameter sets suitable for use differ between the respective decoding methods. The left and right in Fig. 10 are the matrix plots showing the average error matrix $\langle\hat{\boldsymbol{e}}\rangle=\langle\hat{\boldsymbol{x}}\rangle\circ\hat{\boldsymbol{z}}$ over the set of samples $\{\hat{\boldsymbol{x}}\}$ obtained from the MCMC sampler under the parameter sets $A$ and $B$, respectively. The marginal probability $P_{c}$ that $\hat{x}_{ij}$ gives the correct value, i.e., $\hat{x}_{ij}=\hat{z}_{ij}$, is given by $P_{c}=\tfrac{1+\langle{\hat{e}_{ij}}\rangle}{2}$ . So, a negative entry in $\langle\hat{\boldsymbol{e}}\rangle$ indicates it is likely that there is an error in the corresponding entry in $\hat{\boldsymbol{x}}$. We can see that $\langle\hat{\boldsymbol{e}}\rangle$ depends on the chosen parameter sets. The average error matrix $\langle\hat{\boldsymbol{e}}\rangle$ in Fig.10(b) is consistent with the error matrix $\hat{\boldsymbol{e}}$ in Fig.8, because they were sampled with the same parameter set $B$. Note that it is highly likely to fail sampling a code-state by MCMC sampler for the set $B$ (see the lower left plot in Fig.9), while it is highly likely to succeed sampling a code-state by the MCMC sampler for the set $A$. This implies that the state $\hat{\boldsymbol{x}}$ sampled in the first stage of the MCMC-BF decoding is not a code-state, although the state $\hat{\boldsymbol{x}}$ sampled in the MCMC decoding is a code-state. This is quite reasonable because the subsequent BF decoding can correct errors only if $\hat{\boldsymbol{x}}$ sampled at the first stage is a leakage state in the MCMC-BF hybrid decoding. In contrast, the sampled state must be a code-state if it is the correct state in the MCMC decoding. This is why optimal parameter sets $\{\beta_{opt},\gamma_{opt}\}$ differ between the MCMC decoding and MCMC-BF hybrid decoding. It is important to note that this is a consequence of our BF decoding algorithm and is independent of the sampling algorithm in the first stage. Our results suggest that optimal Lagrange weights should be carefully chosen when applying our BF decoding for readouts of QA as they may differ from the optimal parameters for QA when used alone.
@@ V.1 Relationship between MCMC-BF hybrid algorithm and other known BF algorithms
Both the MCMC and BF decoding algorithms are hard-decision algorithms, in which only spin variables are of concern. The MCMC-BF decoding algorithm is considered an extended BF algorithm that incorporates MCMC decoding as a preprocessing step. We discuss the limitations of BF decoding and how they were resolved in the existing extended BF algorithms.
@@
Many researchers have developed sophisticated versions of the weighted BF (WBF) to incorporate reliability information of the hard decisions into the BF algorithm. See details in Refs.[36, 37] , and references therein. Alternatively, there is another approach to incorporate soft information. For example, Wadayama et al. proposed the Gradient Descent Bit Flipping (GDBF) algorithm [28]. They are based on modifications to the inversion function that may improve decoding performance, though this comes at the cost of increased decoding complexity. For example, the inversion functions employed in the BF, WBF, and GDBF algorithms are formally written as
@@
respectively. Here, $\boldsymbol{x}=(x_{1},\ldots,x_{N_{v}})\in\{\pm 1\}^{N_{v}}$ is the current decision for the spin variables (VNs), $\beta$ is a positive real parameter to be adjusted,
@@
is an $i$th syndrome (CN) for the decision $\boldsymbol{x}$ in the spin representation, and $J_{k}$ is identified with channel observation $y_{k}$ in the AWGN channel model. Here, the set $N(i)=\left\{j:H_{ij}=1\right\}$ is the VNs adjacent to an $i$th CN $\left(1\leq i\leq N_{c}\right)$ and the set $M(j)=\left\{i:H_{ij}=1\right\}$ is the CNs adjacent to a $j$th VN $\left(1\leq j\leq N_{v}\right)$. Please refer to Fig.2 for the definition of these sets. Note that as long as $s_{i}\left(\boldsymbol{x}\right)$ is a weight-3 syndrome, $\varDelta_{k}^{\mathrm{(BF)}}(\boldsymbol{x})$ is symmetric for the permutation of the elements of $\boldsymbol{x}$. In other words, $\varDelta_{k}^{\mathrm{(BF)}}(\boldsymbol{x})$ is invariant under exchanging $i\longleftrightarrow j$ for any pair of indices of VN other than $k$, justifying the all-one code-state assumption. In contrast, $\varDelta_{k}^{\mathrm{(BF)}}(\boldsymbol{x})$ is not symmetric if $s_{i}\left(\boldsymbol{x}\right)$ is a weight-4 syndrome. Similarly, because $\varDelta_{k}^{\mathrm{(WBF)}}(\boldsymbol{x})$ involves weight $w_{k}$ in the second term and $\varDelta_{k}^{\mathrm{(GDBF)}}(\boldsymbol{x})$ involves $J_{k}$ in the first term, both depend on $\boldsymbol{J}$. Thus, $\varDelta_{k}^{\mathrm{(WBF)}}(\boldsymbol{x})$ and $\varDelta_{k}^{\mathrm{(GDBF)}}(\boldsymbol{x})$ are not symmetric even if $s_{i}(\boldsymbol{x})$ is a weight-3 syndrome. It is interesting to note that the inversion functions in Eq.(52)-(54) can be formally derived from the following Hamiltonians:
@@
where $X=\mathrm{BF}$, $\mathrm{WBF}$, or $\mathrm{GDBF}$. It follows that if $\varDelta_{k}^{(X)}(\boldsymbol{x})<0$, flipping the $k$th spin reduces the total energy of the spins. Therefore, the BF, WBF, and GDBF decoding algorithms can be considered deterministic algorithms that determine the most suitable spins to be flipped to reduce the energy $H_{k}^{(X)}(\boldsymbol{x})$ based on the inverse function $\varDelta_{k}^{(X)}(\boldsymbol{x})$. In addition, in contrast to the BF algorithm, the BP algorithm essentially considers soft information, as it calculates a consistent marginal probability $P\left(x_{i}|\boldsymbol{y}\right)$ by exchanging real-valued messages between VNs and CNs. On the other hand, the inversion function and the associated Hamiltonian for the MCMC decoding are formally written by
@@
In addition, MCMC decoding in the first stage also solves another limitation of the BF algorithm in the second stage. The MCMC sampling is stochastic in contrast to our deterministic BF algorithm and gradient descent algorithm used in the GDBF algorithm. This stochasticity introduces randomness into the spin-flip selection. It allows spins to flip even when $\varDelta_{k}^{\mathrm{(MCMC)}}(\boldsymbol{x})>0$, which provides an escape from spurious local minima and makes it more likely to arrive at the neighborhood of the global minimum of $H^{\mathrm{(MCMC)}}(\boldsymbol{x})$. In this way, the stochastic spin-flip selection helps find the global minimum of $H^{\mathrm{(MCMC)}}(\boldsymbol{x})$. A similar stochasticity can be incorporated directly into the GDBF algorithm, either by adding a noise term to the inversion function $\varDelta_{k}^{\mathrm{(GDBF)}}(\boldsymbol{x})$ (Noisy GDBF [29]) or by taking stochastic spin-flip selection into account in the algorithm (Probabilistic GDBF [38]). Briefly speaking, our MCMC-BF hybrid decoding algorithm can be considered an alternative to the BP algorithm and other existing BF decoding algorithms. Our simulation suggests that it is important to properly control the fluctuations of $\boldsymbol{x}$ during MCMC sampling by adjusting the two Lagrange weights $\left\{\beta,\gamma\right\}$ to optimize MCMC-BF decoding. It also suggests that the contribution of the penalty term to the Hamiltonian must be small enough not to force $\boldsymbol{x}$ into a valid code-state. Controlling Lagrange weights is important for switching between the two decoding modes. Fig.11 shows schematic diagrams illustrating two decoding modes: (a) MCMC decoding and (b) MCMC-BF hybrid decoding, respectively. Decoding is formulated as a constrained COP. The large and small ellipses indicate the search space and the code-state space that minimizes the penalty term in the Hamiltonian. In MCMC decoding, the MCMC samples both the code-states (feasible solution) and leakage states (infeasible solution) (Fig.11(a)). In some occasions, the correct code-state (solution) may be sampled. In MCMC-BF hybrid decoding, on the other hand, the first-stage MCMC sampler samples only the leakage state, and the second-stage BF decoding maps the leakage state to the correct code-state occasionally (Fig.11(b)). Our simulations showed that the MCMC-BF hybrid decoding (Fig.11(b)) was over two orders of magnitude more efficient than the MCMC decoding (Fig.11(a)) if we ignore the decoding cost of the post-readout BF decoding and compare performance simply by the number of required MCMC samples needed to obtain at least one correct code-state.